# StyleCap: Automatic Speaking-Style Captioning from Speech Based on Speech and Language Self-supervised Learning Models

### 28 Nov 2023 v1

@@ Abstract
We propose StyleCap, a method to generate natural language descriptions of speaking styles appearing in speech. Although most of conventional techniques for para-/non-linguistic information recognition focus on the category classification or the intensity estimation of pre-defined labels, they cannot provide the reasoning of the recognition result in an interpretable manner. As a first step towards an end-to-end method for generating speaking-style prompts from speech, i.e., automatic speaking-style captioning, StyleCap uses paired data of speech and natural language descriptions to train neural networks that predict prefix vectors fed into a large language model (LLM)-based text decoder from a speech representation vector. We explore an appropriate text decoder and speech feature representation suitable for this new task. The experimental results demonstrate that our StyleCap leveraging richer LLMs for the text decoder, speech self-supervised learning (SSL) features, and sentence rephrasing augmentation improves the accuracy and diversity of generated speaking-style captions. Samples of speaking-style captions generated by our StyleCap are publicly available111https://ntt-hilab-gensp.github.io/icassp2024stylecap/.
@@ 1 Introduction
Another crucial factor for better para-/non-linguistic information recognition is the explainability of the recognition results. Although some applications such as healthcare will require not only the recognition result but also its reasoning [9], which should be interpretable for users, most existing techniques focus on the category classification or the intensity estimation of labels defined by authors or dataset developers. One possible approach to solve this issue would be representing the evidence, i.e., para-/non-linguistic information in speech, with a natural language description. Several studies have attempted to acquire the relationship between para-/non-linguistic information and natural language descriptions [10, 11]. However, these methods still output class labels such as emotional states.
From a different viewpoint, generating natural language descriptions from input audio/image data can be regarded as a captioning task. For instance, audio captioning [12] and image captioning [13] are tasks to describe content information (objects and its behavior) in audio and image data, respectively. In these tasks, the remarkable progress of large-scale pre-trained deep neural network (DNN) models [14, 15, 16], has achieved the audio/image captioning performance better than conventional methods. Unlike these tasks, this paper focuses on para-/non-linguistic information rather than content (linguistic) information in speech. However, since this information cannot be directly observed in input speech [17], generating captions of such information can be regarded as a new challenge.
@@ 3.2 Experimental conditions
As for the speech encoder, we used three types of speech feature representations: mel-spectrogram, speaker embeddings for speaker verification, and hidden vectors of a speech SSL model for the performance comparison. We extracted 80-dimensional mel-spectrograms from speech waveforms with 10 ms frame shift. We also used WavLM BASE+444https://huggingface.co/microsoft/wavlm-base-plus as the speech SSL model and extracted weighted-sum representation for each frame. To obtain a fixed-length vector representation for the mel-spectrogram and speech SSL model, a stack of 4-layer BLSTMs and MHA with 8 heads was used as an aggregation module. As for the speaker embeddings, 512-dimensional x-vectors [23] obtained from the pre-trained WavLM BASE+ for speaker verification555https://huggingface.co/microsoft/wavlm-base-plus-sv was used. As for the text decoder, we used two types of pre-trained LLMs: GPT-2666https://huggingface.co/gpt2 [24] and Llama 2 (7B)777https://huggingface.co/meta-llama/Llama-2-7b [19]. The former is the same setting as ClipCap [18], while the latter can be regarded as the richer one. The numbers of model parameters for GPT-2 and Llama 2 were 125M and 7B, respectively. The dimensions of word embeddings for each model were 768 and 4,096, respectively. The mapping network consisted of 8-layer Transformer-encoders. The prefix length and dropout ratio were set to 40 and 0.2, respectively. These parameters were empirically determined. During the model training, we only trained the mapping network, the aggregation module in the speech encoder, and the weight coefficient for each layer in WavLM (i.e., ) used in the SSL feature aggregation process. The model parameters of LLMs and WavLM were frozen. All modules are trained in an end-to-end manner using the cross entropy loss between the generated captions and the ground-truth captions. Each model was trained without sentence rephrasing augmentation at 20 epochs or with the augmentation at 10 epochs because the augmentation doubles the data size. The batch size was set to 16.

### 27 Dec 2023 v2

@@ Abstract
We propose StyleCap, a method to generate natural language descriptions of speaking styles appearing in speech. Although most of conventional techniques for para-/non-linguistic information recognition focus on the category classification or the intensity estimation of pre-defined labels, they cannot provide the reasoning of the recognition result in an interpretable manner. StyleCap is a first step towards an end-to-end method for generating speaking-style prompts from speech, i.e., automatic speaking-style captioning. StyleCap is trained with paired data of speech and natural language descriptions. We train neural networks that convert a speech representation vector into prefix vectors that are fed into a large language model (LLM)-based text decoder. We explore an appropriate text decoder and speech feature representation suitable for this new task. The experimental results demonstrate that our StyleCap leveraging richer LLMs for the text decoder, speech self-supervised learning (SSL) features, and sentence rephrasing augmentation improves the accuracy and diversity of generated speaking-style captions. Samples of speaking-style captions generated by our StyleCap are publicly available111https://ntt-hilab-gensp.github.io/icassp2024stylecap/.
@@ 1 Introduction
Another crucial factor for better para-/non-linguistic information recognition is the explainability of the recognition results. Some applications such as healthcare will require not only the recognition result but also its reasoning [9], which should be interpretable for users. However, most existing techniques focus on the category classification or the intensity estimation of labels defined by authors or dataset developers. One possible approach to solve this issue would be representing the evidence, i.e., para-/non-linguistic information in speech, with a natural language description. Several studies have attempted to acquire the relationship between para-/non-linguistic information and natural language descriptions [10, 11]. However, these methods still output class labels such as emotional states.
From a different viewpoint, generating natural language descriptions from input audio/image data can be regarded as a captioning task. For instance, audio captioning [12] and image captioning [13] are tasks to describe content information (objects and its behavior) in audio and image data, respectively. In these tasks, the remarkable progress of large-scale pre-trained deep neural network (DNN) models [14, 15, 16], has achieved the audio/image captioning performance better than conventional methods. Unlike these tasks, this paper focuses on para-/non-linguistic information rather than content information in speech. However, since this information cannot be directly observed in input speech [17], generating captions of such information can be regarded as a new challenge.
@@ 3.2 Experimental conditions
As for the speech encoder, we used three types of speech feature representations: mel-spectrogram, speaker embeddings for speaker verification, and hidden vectors of a speech SSL model for the performance comparison. We extracted 80-dimensional mel-spectrograms from speech waveforms with 10 ms frame shift. We also used WavLM BASE+444https://huggingface.co/microsoft/wavlm-base-plus [23] as the speech SSL model and extracted weighted-sum representation for each frame. To obtain a fixed-length vector representation for the mel-spectrogram and speech SSL model, a stack of 4-layer BLSTMs and MHA with 8 heads was used as an aggregation module. As for the speaker embeddings, 512-dimensional x-vectors [24] obtained from the pre-trained WavLM BASE+ for speaker verification555https://huggingface.co/microsoft/wavlm-base-plus-sv was used. As for the text decoder, we used two types of pre-trained LLMs: GPT-2666https://huggingface.co/gpt2 [25] and Llama 2 (7B)777https://huggingface.co/meta-llama/Llama-2-7b [19]. The former is the same setting as ClipCap [18], while the latter can be regarded as the richer one. The numbers of model parameters for GPT-2 and Llama 2 were 125M and 7B, respectively. The dimensions of word embeddings for each model were 768 and 4,096, respectively. The mapping network consisted of 8-layer Transformer-encoders. The prefix length and dropout ratio were set to 40 and 0.2, respectively. These parameters were empirically determined. During the model training, we only trained the mapping network, the aggregation module in the speech encoder, and the weight coefficient for each layer in WavLM (i.e., ) used in the SSL feature aggregation process. The model parameters of LLMs and WavLM were frozen. All modules are trained in an end-to-end manner using the cross entropy loss between the generated captions and the ground-truth captions. Each model was trained without sentence rephrasing augmentation at 20 epochs or with the augmentation at 10 epochs because the augmentation doubles the data size. The batch size was set to 16.
@@ 4 Conclusions
Acknowledgements: This work was supported by JST, ACT-X Grant Number JPMJAX23CB, Japan.